OpenStudy (anonymous):

Solve the following equations for x. Identify any extraneous solutions. Show all work.

OpenStudy (anonymous):



OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

You'll want to isolate that radical expression.

OpenStudy (anonymous):

Bring that 4 into the right side: \[\sqrt{x+2}=x-4\] then square both sides: \[x+2=(x-4)^2\] Expand the right side: \[x+2=x^2-8x+16\] Express as a quadratic equation: \[x^2-9x+14=0\] Factor out. The roots are are now obvious. \[(x-7)(x-2)\] Try substituting 7 & 2 to the original equation and see if both sides are equal.

OpenStudy (anonymous):

2 is an extraneous root. Verify it.